if 2 is subtracted from four times a number, the result is 3 more than six times the number. find the number
Answer by ewatrrr(24785)


Hi,
Let x represent the number
Write as we Read
4x - 2 = 6x + 3
-5 = 2x
-5/2 = x
checking our answer
4*-5/2 - 2 = -12 = 6*-5/2 + 3